OpenStudy (anonymous):

To order tickets for a show, there is a $5.00 service fee, plus the cost of the tickets. If the cost of the whole order is $105.00, and 8 tickets were ordered, how much does each ticket cost?

OpenStudy (anonymous):

f(x) = 5 + 8x = 105

OpenStudy (anonymous):

plug and chug

OpenStudy (anonymous):

5 + 8x = 105 8x = 100 x = 12.5